OpenStudy (anonymous):

The distance (y), in miles, of two cars from their starting points at certain times (x), in hours, are shown by equations below: CAR A: y = 60x + 10 CAR B: y = 40x + 70 After how many hours will the two cars be at the same distance from their starting point and what will that distance be? A. 2 hours, 150 miles B. 2 hours, 190 miles C. 3 hours, 150 miles D. 3 hours, 190 miles

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

let's first solve for x \(60x+10=40x+70\) subtract 10 \(60x=40x+60\) subtract 40x \(20x=60\) divide by 20 \(x=3\) so option A and B are out

Mehek (mehek14):

now we solve for y by putting 3 for x in the equation

Mehek (mehek14):

\(y=60*3+10\\ 60*3=180\\ 180+10=190\\ y=190\)

Mehek (mehek14):

so D

Mehek (mehek14):

@itsmee

OpenStudy (anonymous):

ohh okaay thanksss sooo muchh